The marginal cost of a trade has been found to be MC=3Q^2+8Q+400.



determine the total virable cost of producing 100 units of the trades product. And show by graph

1
Expert's answer
2022-01-02T18:20:47-0500

Solution:

MC = 3Q2 + 8Q + 400

First, derive total cost:

Now to derive TC from it we have to integrate over the MC equation. Therefore, we will get:

TC= Q3 + 4Q2 + 400Q

TC = TVC + TFC

Since there is no fixed cost, we assume the above is the Total Variable Cost (TVC):

TVC = Q3 + 4Q2 + 400Q

TVC for producing 100 units:

Q = 100

TVC = 1003 + 4(1002) + 400(100) = 1,000,000 + 40,000 + 40,000 = 1,080,000

TVC for producing 100 units = 1,080,000